A ball is released from a tower at a height of 100 meters toward the roof of another tower that is 25 meters high. The horizontal distance between the two towers is 20 meters. With what horizontal velocity should the ball be imparted so that it lands on the rooftop of the second building?
Answer #1
h = Yo*t + 0.5g*t^2 = 100 - 25 = 75 m.
o + 4.9t^2 = 75.
t^2 = 15.31 m.
Tf = 3.91 s. = Time to fall to roof.

Dx = Xo * Tf = 20 m.
Xo * 3.91 = 20.
Xo = 20 / 3.91 = 5.11 m/s. = hor. velocity.
answered by: Eden